OpenStudy (anonymous):

Can someone show me how to do this problem im stuck on it!!! What is the x-coordinate of the x-intercept of the line that passes through points (-5, 3) and (1, -3)?

OpenStudy (anonymous):

@dan815 @chmvijay @iribeiro1602

OpenStudy (anonymous):

Slope of the line is -1. You know how to get that using (y2-y1)/(x2-x1)?

OpenStudy (anonymous):

Now use the point slope formula: y-y1=m(x-x1) where m = -1 x1, y1 is any point on your line. Lets use (1,-3). So x1=1, y1=-3

OpenStudy (anonymous):

y+3 = -1(x-1) y+3 = -x +1 To get the x int set y=0 0+3 = -x + 1 x = 1-3 x = -2 Answer is -2

OpenStudy (anonymous):

If you want me to go back over any step here let me know.

OpenStudy (anonymous):

oh i see what i was doing wrong haha LMFAO

OpenStudy (anonymous):

thank you big time lol i would of been stuck there all day trying to remember the process

OpenStudy (anonymous):

sure. no problem